OpenStudy (anonymous):

Shirley wants to buy a skateboard for $64. She has $98 in her account. She spent $10.85 to buy stationary. She also wants to buy some cookies for $1.65 each. What is the maximum number of cookies, n, that Shirley can buy so that she has enough money left to buy the skateboard?

OpenStudy (abhisar):

Hey there! According to the question we can set up an equation to find out the number of cookies n she can buy. Things she have or must buy are: \(\bullet\) $64 Skater board \(\bullet\) $ 10.85 stationary \(\bullet \) cookies worth $ 1.65 each So, if she bought n number of cookies then she actually spent \(\displaystyle n \times 1.65\) $ on cookies.

OpenStudy (abhisar):

So, finally we can set up an equation as follows: \(\tt 64+10.85+(1.65n)=98\) , solve this equation and find the value of n

OpenStudy (anonymous):

ok

OpenStudy (abhisar):

Can you do that?

OpenStudy (abhisar):

Also post mathematics questions inn mathematics section. http://openstudy.com/study#/groups/mathematics

OpenStudy (anonymous):

-14